# Natural solution of SUSY $\mu$ problem from modulus stabilization in modular flavor model

### v3

@@ 4 Stabilization of modulus field with nilpotent Goldstino ${\textbf{X}}_{nl}$ superfield
      for values of $\tau$ within the fundamental region. To obtain a vanishing (very tiny positive) vacuum energy $\Lambda^{4}\sim 10^{-120}M_{Pl}^{4}\approx 0$, we can find the parameter points that are solutions of ${H(\tau)}=0$ and also satisfy the minimum condition (with respect to $\tau$ and $\tau^{*}$) with $\tilde{W}_{M}=0$. It can be checked that solutions of ${H(\tau)}=0$ can also satisfy the minimum condition $\tilde{W}_{M}=0$, which will be clear shortly. Therefore, the stabilized modulus VEV is located at the solutions of ${H(\tau)}=0$ with (almost) vanishing vacuum energy in this case.
    We now focus on the general case with $f_{0},c_{1},X_{nl}\neq 0$. It can be calculated that the extreme condition for $\theta$ is
    $\frac{\partial}{\partial\theta}V=-\frac{2f_{0}c_{1}^{3}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\rho\sin\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]=0.$ (47)
    First, we should note that, as $[(2\Im\tau)|\eta(\tau)|^{4}]^{-1}$ diverges when $\Im\tau=\infty$, the $\tau=i\infty$ point is not a solution of the extreme condition for $\theta$.
    For generic $\theta$ (with $\theta\neq 0,\pi$ to render $X_{nl}$ not real) in the case $f_{0},c_{1},\rho\neq 0$, the extreme condition for $\theta$ reduces to the constraint
    $\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0~{}.$ (48)
    After substituting this constraint in the extreme condition for $\rho$, we have
    $\frac{\partial}{\partial\rho}V=4\rho f_{0}^{2}\frac{\left|H(\tau)\right|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}=0~{}.$ (49)
    Such an extreme condition for $\rho$ can be satisfied with
    $\left|H(\tau)\right|^{2}=0~{}.$ (50)
    Thus, we have the extreme conditions for the case with $f_{0},c_{1}\neq 0$ and a non-vanishing non-real $X_{nl}$
    $\left|\tilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0~{}.$ (51)
    Substituting back into the scalar potential, we can see that the potential energy at the minimum vanishes. Therefore, the presence of a non-trivial $X_{nl}$ sector, whose contributions can uplift the scalar potential, can naturally lead to a flat Minkowski universe in this case.
    From the expression of $H(\tau)$, the constraint $H(\tau)=0$ requires
    $j(\tau)^{n/3}[j(\tau)-1728]^{m/2}=0,~{}~{}{\rm or}~{}~{}~{}~{}{\cal P}(j(\tau))=0.$ (52)
    Obviously, $\tau=i,\omega$ are its solutions for $m\geq 1$ and $n\geq 1$, respectively. As ${\cal P}(j(\tau))$ is an arbitrary polynomial function of $j(\tau)$, from the fundamental theorem of algebra, any complex value of $j(\tau)$ can be the root of typical chosen polynomial equation ${\cal P}(j(\tau))=0$, leading to no interesting constraint at all unless we specify the concrete form of ${\cal P}(j(\tau))$.
@@
    $\tilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0~{},$ (53)
    which satisfies $H(\tau)=0$. Such a constraint reduces to
    $H^{\prime}(\tau)=\left\{j(\tau)^{n/3}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))\right\}^{\prime}=0~{},$ (54)
    $\Rightarrow$ $j^{\prime}(\tau)\left[\frac{n}{3}j(\tau)^{n/3-1}[j(\tau)-1728]^{m/2}{\cal P}(j(\tau))+\frac{m}{2}j(\tau)^{n/3}[j(\tau)-1728]^{m/2-1}{\cal P}(j(\tau))\right.$
    $\left.+j(\tau)^{n/3}[j(\tau)-1728]^{m/2}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$
    $\Rightarrow$ $H(\tau)j^{\prime}(\tau)\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{}.$
    So $H^{\prime}(\tau)=0$ can trivially hold when $H(\tau)=0$.
    On the other hand, $H^{\prime}(\tau)=0$ can also be satisfied by the other two conditions
    $\left[\frac{n}{3}\frac{1}{j(\tau)}+\frac{m}{2}\frac{1}{j(\tau)-1728}+\frac{1}{{\cal P}(j(\tau))}\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}\right]=0~{},$ (55)
    or $j^{\prime}(\tau)=0$.
    We have the following discussions for the condition (55)
    • $m+n\neq 0$:
      When $n\neq 0$ and $m=0$, condition (55) reduces to
      $\frac{d{\cal P}(j(\tau))}{{\cal P}(j(\tau))}=-\frac{n}{3}\frac{dj(\tau)}{j(\tau)}~{},$ (56)
      which gives
      ${\cal P}(j(\tau))=C_{0}\left[j(\tau)\right]^{-n/3}~{}.$ (57)
      Unless $n\in 3\mathbb{Z}$ is negative, ${\cal P}(j(\tau))$ is not a polynomial of $j(\tau)$. Similarly, when $m\neq 0$ and $n=0$, we can deduce that $m\in 2\mathbb{Z}$ should be negative. For $m\neq 0$ and $n\neq 0$, similar deduction can give
      ${\cal P}(j(\tau))=C_{0}^{\prime}\left[j(\tau)\right]^{-n/3}\left[{j(\tau)-1728}\right]^{-m/2}~{},$ (58)
      which also requires both $m\in 2\mathbb{Z}$ and $n\in 3\mathbb{Z}$ being negative to guarantee ${\cal P}(j(\tau))$ being a polynomial of $j(\tau)$. As $m,n$ are defined to be positive integers, this condition requires $m=n=0$.
    • $m=0$ and $n=0$:
      With the constraints $m=n=0$, the condition (55) reduces to
      $\frac{\partial{\cal P}(j(\tau))}{\partial j(\tau)}=0~{},$ (59)
      which gives
      ${\cal P}(j(\tau))={\rm Constant}.$ (60)
      As $H(\tau)=0$, this constant should be zero.
    Therefore, the condition (55) requires a vanishing $H(\tau)$. For the condition $j^{\prime}(\tau)=0$, from the derivative of the Klein $j$-function
    $\frac{\partial}{\partial\tau}j(\tau)=-2\pi i\frac{{E_{6}(\tau)}{E_{4}^{2}(\tau)}}{[\eta(\tau)]^{24}}=0~{},$ (61)
    we can see that its solutions are $\omega$ and $i$, leading to no new additional constraints. Therefore, the extreme conditions (51) can reduce to the requirement of vanishing $H(\tau)$.
    Before we conclude this scenario, it should be noted that $\theta=0,\pi$, which corresponds to real $X_{nl}=\pm\rho$ with positive $\rho$, are also solutions of the the extreme condition (47) for $\theta$. In this case, the scalar potential reduces to
@@
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right.$
    $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}\rho^{2}\pm 2f_{0}c_{1}^{3}\rho+\rho^{2}c_{1}^{6}\frac{}{}\right)~{}$
    -- $\left.3\left|{H(\tau)}\right|^{2}\left(f_{0}^{2}\rho^{2}+c_{1}^{6}\pm 2f_{0}\rho c_{1}^{3}\right)\right\}~{}$
    $+$ $\frac{\rho^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|{H(\tau)}\right|^{2}\right\}~{}.$
    The scalar potential, which is a quadratic polynomial of $\rho$, can be written as
    $V=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left[C_{1}\rho^{2}\pm C_{2}\rho+C_{3}\right]~{},$ (63)
@@
    The scalar potential has a minimum with the corresponding $\rho$ parameter
    $\rho_{0}=\mp\frac{C_{2}}{2C_{1}}~{},$ (65)
    and the potential energy
    $\left.V\right|_{\rho=\rho_{0}}=C_{3}-\frac{C_{2}^{2}}{2C_{1}}~{}.$ (66)
    The modulus VEV is determined by the minimum conditions for the potential (66) with respect to $\tau,\tau^{*}$, and the condition $C_{2}^{2}=2C_{1}C_{3}$ should be established to obtain a vanishing (almost vanishing tiny positive) vacuum energy. It can be checked from previous discussions that solutions of ${H(\tau)}=0$ can also satisfy such minimum conditions.
    Thus, we can conclude that, in the scenario with $k_{X}=0$ for $f_{0}\neq 0$, the extrema of the scalar potential are located at the points determined by $H(\tau)=0$, and the corresponding vacuum is of Minkowski type. As noted previously, the constraint $H(\tau)=0$ predicts that the fixed points $z_{0}=\omega,i$ are always the local extrema unless $m=n=0$. Other solutions of $\tau$ from ${\cal P}(j(\tau))=0$ are essentially possible because of the arbitrariness of the polynomial form of ${\cal P}(j(\tau))$. Therefore, other minima, which are determined by (44) to lie at the boundary of the fundamental domain or at the imaginary axis, can also be possible.
    In fig.1, we show the energies of the scalar potential at modulus values taken within the fundamental domain with various parameter choices. It can be seen that the potential energies are non-negative when $f_{0}\neq 0$.
@@
    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ for various parameter choices when $f_{0}\neq 0$. We adopt $\mathcal{P}(j)=1$ for the left panel and $\mathcal{P}(j)=j-4.5764069879884376\times 10^{-11}$ for the right panel.
  • Scenario with small $X_{nl}$ (when $|X_{nl}|\rightarrow 0$):
    In this case, at small $|X_{nl}|$ region, the $|X_{nl}|^{2}$ terms can be neglected in addition to the constraint $X_{nl}^{2}=0$. Note that the $|X_{nl}|^{2}$ terms will dominate the scalar potential at very large $|X_{nl}|$ values, ensuring that the potential can be bounded from below.
    The scalar potential can be simplified to
    $V(X_{nl},{X}^{*}_{nl},\tau,{\tau}^{*})$ (67)
    $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\left[\eta(\tau)\right]^{2k_{X}}\right]\right)\right.$
    $+$ $\left.\frac{(2\Im\tau)^{2}}{3}\left(2\Re\left[k_{X}f_{0}X_{nl}\left[\eta(\tau)\right]^{2k_{X}}\tilde{W}_{E}\right]\right)H(\tau)c_{1}^{3}\tilde{W}_{M}^{*}\right]$
    $+$ $(2\Im\tau)^{k_{X}}\left|H(\tau)\right|^{2}\left(f_{0}^{2}\left|\eta(\tau)\right|^{4k_{X}}+2\Re\left[f_{0}\eta(\tau)^{2k_{X}}\frac{{X}_{nl}c_{1}^{3}}{(2\Im\tau)^{k_{X}}}\right]\right)$
    $+$ $2\Re\left[\frac{i(2\Im\tau)k_{X}X_{nl}}{3}\left(f_{0}[\eta(\tau)]^{2k_{X}}H(\tau)\right)^{*}\tilde{W}_{M}c_{1}^{3}\right]$
    -- $\left.3\left(c_{1}^{6}+2\Re\left[f_{0}X_{nl}c_{1}^{3}\right]\right)\left|{H(\tau)}\right|^{2}\frac{}{}\right\}~{},$
    with
    $\tilde{W}_{E}$$\equiv$ $\frac{i}{2\pi}\hat{G}_{2}(\tau,\tau^{*})+\frac{i}{2\Im\tau}~{}.$ (68)
    It can be seen easily from the linear terms of $X_{nl}$ that the extrema conditions for $X_{nl}$ will not lead to trivial constraints for $\tau$ with arbitrary choices of $k_{X}$. Therefore, other modulus values lying on the boundary of fundamental domain or imaginary axis other than the fixed point, can possibly be the minimum. The initially stabilized modulus VEV without contributions from $X_{nl}$ sector can then be destabilized by new contributions from non-vanishing $X_{nl}$. We will discuss the extrema conditions for the most general case in our subsequent studies.

### v4

@@ 4.1 The case with $k_{X}=0$
    for all values of $\tau$ within the fundamental region. The parameter points for local minimum of the scalar potential should satisfy both $H(\tau)=0$ and $X_{nl}=0$. As $|H(\tau)|^{2}$ can be effectively factored out from $V_{\tau}$ with $P(j(\tau))$ taking the form (86), solutions of $H(\tau)=0$ indeed satisfy $V_{\tau}=0$ in general (except for typical choices of $m,n\leq 2$ at the fixed points). Therefore, in this case, the stabilized modulus VEVs are located at the solutions of $H(\tau)=0$, corresponding to vanishing vacuum energy.
  • Case $f_{0}\neq 0$, $c_{1}\neq 0$:
    We now focus on the general case with $f_{0},c_{1}\neq 0$. It can be calculated that the extremum condition for $\theta$ is
    $\frac{\partial}{\partial\theta}V=-\frac{2f_{0}c_{1}^{3}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\rho\sin\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]=0~,$ (90)
    while the extremum condition for $\rho$ is
    $\frac{\partial}{\partial\rho}V$ $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{2\rho\left[\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}\left(f_{0}^{2}+c_{1}^{6}\right)-2c_{1}^{6}\left|H(\tau)\right|^{2}\right]\right$. (91)
    $+$ $\left.2f_{0}c_{1}^{3}\cos\theta\left[\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}\right]\right\}=0~.$
    It should be noted that the point $\tau=i\infty$ does not satisfy the extremum condition for $\theta$ since $(2\Im\tau)|\eta(\tau)|^{4}]^{-1}$ diverges as $\Im\tau\to\infty$.
    As discussed previously, the $|H(\tau)|^{2}$ factor can be extracted from $\left|\widetilde{W}_{M}\right|^{2}$ for non-vanishing $H(\tau)$. Besides, $\left|\widetilde{W}_{M}\right|^{2}$ vanishes when $H(\tau)=0$ for choices of $P(j(\tau))$ in the form (86) and $m,n\neq 1$. Therefore, $|H(\tau)|^{2}=0$, which in general satisfies the extremum condition $V_{\tau}=0$ for $\tau$, can also in general satisfy the extremum conditions $V_{\rho}=V_{\theta}=0$ for $X_{nl}$. So we can conclude that modulus VEVs that satisfy $|H(\tau)|^{2}=0$ for any values of $X_{nl}$ are always local extrema of the scalar potential.
    The extremum conditions for $\theta$ and $\rho$ can also be satisfied in the following cases:
    • From (90), for $\sin\theta\neq 0$ (that corresponds to a non-real $X_{nl}$ with $\theta\neq 0,\pi$), we can get the condition for modulus $\tau$
      $\frac{(2\Im\tau)^{2}}{3}\left|\widetilde{W}_{M}\right|^{2}-2\left|H(\tau)\right|^{2}=0,$ (92)
      which is independent of the $\rho$ value. After substituting this condition into (91), we can obtain the extremum condition for the $\rho$ parameter:
      $\frac{\partial}{\partial\rho}V=4\rho f_{0}^{2}\frac{\left|H(\tau)\right|^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}=0.$ (93)
      Such an extremum condition for $\rho$ can be satisfied with
      $\left|H(\tau)\right|^{2}=0.$ (94)
      Thus, we have the extremum conditions for the case with $f_{0},c_{1}\neq 0$ and a non-vanishing non-real $X_{nl}$ value:
      $\left|\widetilde{W}_{M}\right|^{2}=\left|H(\tau)\right|^{2}=0.$ (95)
      Substituting back into the scalar potential, we can see that the potential energy at this extremum vanishes.
@@
      $\widetilde{W}_{M}\equiv H^{\prime}(\tau)-\frac{3iH(\tau)}{2\pi}\hat{G}_{2}(\tau,\tau^{*})=0,$ (96)
      which satisfies $H(\tau)=0$. Discussions on the solutions of $H^{\prime}(\tau)=0$ are given in Appendix C. For the fixed points $\tau=\omega,i$, they are always the local extrema of the scalar potential when $m,n>1$. When $m<1$ or $n<1$, whether the fixed points $\tau=\omega,i$ are local extrema can easily be figured out case by case, with the relevant results agreeing with the single modulus case in Leedom:2022zdm . The extrema conditions (95) and (66) of the scalar potential can also be satisfied for non-fixed points solutions of $H(\tau)=0$ when $P(j(\tau))$ takes the form (72), and the corresponding vacua are always of Minkowski type. The Minkowski vacua determined by $H(\tau)=0$ are independent of the specific value of $X_{nl}$. To identify the global minimum of the scalar potential, it is necessary to evaluate the potential energies for all extrema, including the case where $X_{nl}$ is real.
    • Real values of $X_{nl}$ with $\theta=0$ or $\theta=\pi$.
      In this case, the scalar potential reduces to
@@
      $=$ $\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}\left(f_{0}^{2}\tilde{X}^{2}+c_{1}^{6}+2f_{0}\tilde{X}c_{1}^{3}\right)\right$.
      $+$ $\left|H(\tau)\right|^{2}\left(f^{2}_{0}+2f_{0}^{2}\tilde{X}^{2}+2f_{0}c_{1}^{3}\tilde{X}+\tilde{X}^{2}c_{1}^{6}\right)$
      - $\left.3\left|H(\tau)\right|^{2}\left(f_{0}^{2}\tilde{X}^{2}+c_{1}^{6}+2f_{0}\tilde{X}c_{1}^{3}\right)\right\}$
      $+$ $\frac{\tilde{X}^{2}}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left\{\frac{(2\Im\tau)^{2}}{3}\left|\tilde{W}_{M}\right|^{2}c_{1}^{6}+\left|H(\tau)\right|^{2}f^{2}_{0}-3c_{1}^{6}\left|H(\tau)\right|^{2}\right\},$
      for real values of $X_{nl}$, which we denote simply as $\tilde{X}$. It is a quadratic polynomial of $\tilde{X}$, which can be rewritten as
      $V(\tilde{X},\tau,\tau^{*})=\frac{1}{(2\Im\tau)^{3}|\eta(\tau)|^{12}}\left[C_{1}\tilde{X}^{2}+C_{2}\tilde{X}+C_{3}\right]~,$ (98)
@@
      Obviously, the scalar potential is real for all (real) values of $\widetilde{X}$. The scalar potential is bounded from below for $\widetilde{X}\in\mathbb{R}$ when the coefficients of the quadratic polynomial in (98) satisfy $C_{1}>0$ (and $C_{2}=0$ when $C_{1}=0$, rendering the scalar potential finite).
@@
      For $C_{1}>0$, the scalar potential has a minimum with the corresponding $\widetilde{X}$ value and the potential energy
      $\left.V\right|_{min;\tilde{X}=\frac{-C_{2}}{2C_{1}}}$ $=$ $C_{3}-\frac{C_{2}^{2}}{4C_{1}}~.$~ (100)
      To determine the extrema of $\left.V\right|_{min}$, we need to determine
      $\frac{\partial}{\partial\tau}\left[C_{3}-\frac{C_{2}^{2}}{4C_{1}}\right]=\frac{\partial}{\partial\tau^{*}}\left[C_{3}-\frac{C_{2}^{2}}{4C_{1}}\right]=0~.$ (101)
      These extremum conditions, after factoring out a common non-vanishing $|H(\tau)|^{2}$ term, yield non-trivial equations for $\tau$ that generally require numerical solution.
      For $C_{1}=0$, it can be checked that $C_{2}=0$ when stationary condition $H(\tau)=0$ for modulus VEV $\tau$ holds, which corresponds to Minkowski vacua for all values of $\tilde{X}$. Therefore, in line with earlier discussions, VEVs of modulus field that satisfy $H(\tau)=0$ for all values of ${X}_{nl}$ (whether real or complex) represent local minima of the scalar potential.
      The pseudo-modulus $X_{nl}$, i.e. a continuous space of vacua with degenerate tree-level vacuum energies, can be lifted by SUSY breaking effects. Taking into account soft SUSY breaking mass term $m_{X}^{2}|X_{nl}|^{2}$ from additional SUSY breaking sources with proper mediation mechanism (or from the one-loop Coleman-Weinberg potential of $X_{nl}$ after including its couplings to matter fields, for example, of the O’Raifeartaigh type $\lambda_{ij}X_{nl}+M_{ij}]\phi_{i}\phi_{j}$ in the superpotential), the scalar potential for $X_{nl}$ taking the form $V(X_{nl})\approx|\tilde{C}|^{2}f_{0}^{2}+m_{X}^{2}|X_{nl}|^{2}$ can stabilize the pseudo-modulus at $X_{nl}=0$ in the flat SUSY limit, with $\tilde{C}$ some constant determined by the stabilized value of $\tau$.
[…]
    For the case $C_{1}>0$, the minimum (100) is always located at some value of $\widetilde{X}$ with $\widetilde{X}\ll M_{Pl}$. This minimum can either be the global minimum or merely a local minimum when the bounded-from-below condition is relaxed by allowing $C_{1}\leq 0$.
    We select some benchmark parameter points with various choices of $f_{0},c_{1}$ and calculate numerically the potential energies for all extrema to identify the global minimum. For many benchmark points, the global minimum is determined by the condition $H(\tau)=0$ with vanishing vacuum energy, which are independent of the value of $X_{nl}$  [^2] As noted previously, the pseudo-modulus $X_{nl}$ will be stabilized at $X_{nl}=0$ after including the SUSY breaking contributions. So we choose the stabilized value $X_{nl}=0$ for the extrema determined by the condition $H(\tau)=0$.. In Fig.1, we show the energies of the scalar potential for modulus values within the fundamental domain and $X_{nl}=0$, considering various parameter choices.
@@
    Figure 1: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=0$ and $X_{nl}=0$ for various parameter choices when $f_{0}\neq 0$ and $f_{0}\gg c_{1}^{2}$. We adopt $\mathcal{P}(j(\tau))=1$ for the left panel and $\mathcal{P}(j(\tau))=(j(\tau)-j(\tau_{0}))^{2}$ with $\tau_{0}=e^{i\frac{2\pi}{3}(1-0.006)}$ for the right panel.
    In the limit $f_{0}\to 0$, this scenario reduces to the one involving only a single modulus field $\tau$, whose vacuum is generally not of Minkowski type and the corresponding vacuum energy can possibly be negative. Therefore, for tiny values of $f_{0}$, the vacuum energy of the global minimum can become negative. Numerical results on the global minima of some benchmark points with non-vanishing vacuum energy are shown in Table 1.
    $m$ $n$ $f_{0}/M_{\mathrm{Pl}}^{2}$ $c_{1}/M_{\mathrm{Pl}}$ $X_{nl}/M_{\mathrm{Pl}}$ $\tau$ $P(j(\tau))$ $V_{min}/M_{\mathrm{Pl}}^{4}$
    0 0 $0.1$ $0.5$ 0.6318-0.6318 $0.0021+1.002i$ 11 0.0525-0.0525
    11 11 $^{-4} 0.10.1$ $0.9994$ $0.19337+0.98275i$ 11 1.25-1.25
    22 0 $10^{-7}$ 0.01-0.01 0.9994-0.9994 $0.4892+0.8733i$ 11 $4.466\times 10^{-5}$
    11 22 $1.128\times 10^{-6}$ 0.01-0.01 0.3322-0.3322 0.1910+0.9830i-0.1910+0.9830i 11 $7.282\times 10^{-5}$
    22 11 $2.974\times 10^{-7}$ 0.01-0.01 $0.2920$ $0.2865+0.9612i$ 11 $7.434\times 10^{-4}$
    11 33 $2.337\times 10^{-6}$ 0.01-0.01 0.4126-0.4126 $0.1588+0.9911i$ 11 $7.103\times 10^{-4}$
    Table 1: The global minima and the corresponding vacuum energies for some benchmark parameter points in the case $k_{X}=0$, $f_{0}\neq 0$ and $c_{1}\neq 0$.
    From a physical perspective, we are particularly interested in the local minimum with an intermediate/small value of $\widetilde{X}$, rather than a global minimum situated at $\widetilde{X}\sim M_{\text{Pl}}$ or $\widetilde{X}\sim-M_{\text{Pl}}$. For typical benchmark parameter choices, an estimation of the lifetime for the local minimum with an intermediate/small value of $\widetilde{X}$ suggests that such a local minimum can be metastable, with a lifetime exceeding the age of the universe. Therefore, we will concentrate on the global (or long-lived meta-stable) minimum with $|X_{nl}|\ll M_{Pl}$ in our subsequent discussions.
    The lifetime of a vacuum can be evaluated semi-classically through a rate of bubble nucleation in unit volume and unit time Coleman ; Coleman1 . The decay probability, which sets the decay rate, is given by $\Gamma/V=A\exp(-B)$, where $B$ is the action of a Euclidean $O(4)$-symmetric bounce solution, which smoothly interpolates between a true-vacuum configuration at its center and the false vacuum whose decay is being described. The prefactor $A$ depends on the functional determinants and represents the loop correction of the fluctuations of the fields. The prefactor $A$ can be approximately given by $A\sim v^{4}$, where $v$ is some characteristic mass scale of the theory under study. We adopt a conservative choice $v\sim{\cal O}(M_{Pl})$ for our estimation of lifetime.
    The generalization to cases involving multiple fields entails determining the bounce configuration by solving a set of coupled ordinary differential equations, subject to the appropriate boundary conditions. For general multi-dimensional field spaces, there will be many bounce solutions with different actions, and the tunneling rate will be dominated by the solution with the lowest action. The field profile, which solves the Euclidean equations of motion, traverses a one-dimensional path in the field space. If one chooses a particular small subset of path in the field space and treats it as a one-field potential, solving the field equations for the one-dimensional potential along this path can give an estimation on the action, which can be close to the actual value of the action 1702-00356 . As the tunneling is expected intuitively to happen through the lowest possible barrier and shortest possible path in the field space, the one with the shortest possible path length which is a straight line between the two minima (or the path with the lowest possible barrier in which a broken line goes through the lowest ridge in the barrier), can act as the one-dimensional approximations of potential along such a path parameterized by the path length in the field space.
    We estimate the decay lifetime with the bounce configuration via the two paths connecting the true and false vacuum for some benchmark parameter choices that can lead to false vacuums at intermediate/small value of $\widetilde{X}$. We estimate the one-dimensional action for each case with the approximation of a triangular potential, whose the bounce action is given approximately Duncan by
    $B\approx\frac{32\pi^{2}}{3}\frac{(\Delta V_{+})^{2}(\Delta\phi)^{4}}{\epsilon^{3}}~.$ (102)
    in the situation where the difference of potential energy between the true and false vacuum $\epsilon$ is very small. Here $\Delta V_{\pm}=V_{Top}-V_{\pm}$ denotes the height for false (and true) vacuum with respect to the top of the barrier and $\Delta\phi=(\phi_{-}-\phi_{+})$ denotes the width of the barrier.
    For a characteristic benchmark point, such as $f_{0}=10^{-7}M^{2}_{Pl}$ and $c_{1}=-0.01M_{Pl}$ with $m=1,n=1$ and $P(j(\tau))=(j(\tau)-j(\tau_{0}))^{2}$ for $\tau_{0}=1.001\times e^{i\frac{2\pi}{3}(1-\frac{1}{36})}$, we find that the false vacuum lies at $(X_{nl},\tau)=(0,1.001\times e^{i\frac{2\pi}{3}\frac{35}{36}})$ with vanishing vacuum energy while the true minimum vacuum lies at $(X_{nl},\tau)=(-0.99999M_{Pl},-0.42825+0.91479i)$ with $V_{min}=-2.49896M_{Pl}^{4}$. The peak of the lowest barrier sits at the parameter point $(X_{nl},\tau)=(-0.46285M_{Pl},-0.40398+0.92017i)$ with $V_{L-peak}=577.87447M_{Pl}^{4}$. We find that the bounce action of the broken line path through the lowest possible barrier leads to smaller action. It can be calculated with (102) that the corresponding bounce action $B\approx 2.28\times 10^{5}$. Therefore, the lifetime of the meta-stable vacuum $\tau=(\Gamma/V)^{-1/4}$ is much larger than the universe. In fact, using the space-time volume of our past lightcone Past:LC , $(VT)_{LC}=0.15/H_{0}^{4}=3.4\times 10^{166}{\rm GeV}^{-4}$ with the Hubble constant $H_{0}=1.44\times 10^{-42}{\rm GeV}$, the probability that we should have seen a bubble by now is $(VT)_{LC}](\Gamma/V)\sim 10^{-9670}$, which is exponentially small.
4.2 The case with $k_{X}=12$
When $k_{X}=12$, interference terms in the scalar potential obscure the analysis of the extrema conditions. Therefore, we leave the most general discussions on this case in our subsequent studies. An important observation is that an $|H(\tau)|^{2}$ factor (for $H(\tau)\neq 0$) can be extracted from the scalar potential. According to the discussions in the case $k_{X}=0$, for any value of $X_{nl}$, in addition to the fixed points $\tau=i,\omega$ for most choices of $m,n$, non-fixed point values of modulus VEVs satisfying $|H(\tau)|^{2}=0$ are local extrema of the scalar potential with vanishing vacuum energy when $P(j(\tau))$ takes the form (72). Other local minima of the scalar potential can be determined numerically from the extrema conditions with respect to $\tau$, $X_{nl}$ and the corresponding Hessian matrix. After specifying numerically all other extrema of the scalar potential, we can determined its global minimum or meta-stable minimum by evaluating the corresponding vacuum energies of various local minima (and the lifetime of the meta-stable minimum). We show in Fig.2 the vacuum energy of the scalar potential for typical benchmark parameter choices with modulus values taken in the fundamental domain in the scenario with $k_{X}=12$.
Figure 2: The values of $V(X_{nl},\bar{X}_{nl},\tau,{\tau}^{*})$ within the fundamental domain in the scenario with $k_{X}=12$ and $X_{nl}=0$ for various parameter choices when $f_{0}\neq 0$ and $f_{0}\gg c_{1}^{2}$. We adopt $\mathcal{P}(j)=1$ for the left panel and $\mathcal{P}(j(\tau))=(j(\tau)-j(\tau_{0}))^{2}$ with $\tau_{0}=1.05\times e^{i\frac{\pi}{2}(1+\frac{1}{6})}$ for the right panel.
In our discussion, we assume the existence of modular symmetry in the SUSY breaking sector. The modular symmetry can be broken by SUSY breaking effects, depending on the transformation property of SUSY breaking source under the modular symmetry, as discussed in Kikuchi:2022pkd . For example, in the moduli mediated SUSY breaking scenario Kikuchi:2022pkd , soft SUSY breaking scalar masses and trilinear terms are invariant under the modular symmetry when the F-term is regarded as a spurion with the modular weight 2-2. In our case, non-vanishing F-term VEV of $X_{nl}$ when $k_{X}\neq 0$ can possibly act as the SUSY breaking source and also break the modular symmetry, whose transformation should be compatible with the modular invariance of $K_{X_{nl}^{\dagger}X_{nl}}F_{X_{nl}}$, possibly leading to a soft SUSY breaking spectrum that is non-invariant under modular symmetry. The form of the low energy soft SUSY breaking spectrum depends on the concrete mediation mechanism of SUSY breaking that mediates the SUSY breaking effect from the hidden sector to the visible sector.